# DP-DyLoRA: Fine-Tuning Transformer-Based Models On-Device under Differentially Private Federated Learning using Dynamic Low-Rank Adaptation

### v2

@@ Abstract
Federated learning (FL) allows clients in an Internet of Things (IoT) system to collaboratively train a global model without sharing their local data with a server. However, clients’ contributions to the server can still leak sensitive information. Differential privacy (DP) addresses such leakage by providing formal privacy guarantees, with mechanisms that add randomness to the clients’ contributions. The randomness makes it infeasible to train large transformer-based models, common in modern IoT systems. In this work, we empirically evaluate the practicality of fine-tuning large scale on-device transformer-based models with differential privacy in a federated learning system. We conduct comprehensive experiments on various system properties for tasks spanning a multitude of domains: speech recognition, computer vision (CV) and natural language understanding (NLU). Our results show that full fine-tuning under differentially private federated learning (DP-FL) generally leads to huge performance degradation which can be alleviated by reducing the dimensionality of contributions through parameter-efficient fine-tuning (PEFT). Our benchmarks of existing DP-PEFT methods show that DP-Low-Rank Adaptation (DP-LoRA) consistently outperforms other methods. An even more promising approach, DyLoRA, which makes the low rank variable, when naively combined with FL would straightforwardly break differential privacy. We therefore propose an adaptation method that can be combined with differential privacy and call it DP-DyLoRA. Finally, we are able to reduce the accuracy degradation and word error rate (WER) increase due to DP to less than 2% and 7% respectively with 1 million clients and a stringent privacy budget of $\bm{\epsilon}=2$.
@@ I Introduction
Today, transformer-based models [1] are becoming increasingly common in Internet of Things (IoT) systems for a wide range of applications such as natural language understanding (NLU), automatic speech recognition (ASR) and image classification [2, 3]. Compared to models such as recurrent neural networks (RNNs) and convolutional neural networks (CNNs), transformer-based models are known to have several advantages such as being better at handling long-range input dependencies and more efficient for training and inference due to parallel input processing [1]. Pre-training and fine-tuning transformers is the dominant approach for building models with state-of-the-art performance [4, 5]. These models are particularly suitable for deployment in IoT systems since they can be pre-trained on massive unlabelled data at the central server without much human effort, and only a small amount of data is required per client when fine-tuning in collaboration with other IoT clients for downstream tasks.
@@
Comprehensive experiments for DP-PEFT methods are missing from the literature. Experiments in existing works [19, 20, 21] fail to address realistic system properties such as a massive number (millions) of clients in an IoT system [22]. Works such as [20] and [21] show experiments for only a single domain and a single type of DP-PEFT method. Only speech is considered in [19], but it only evaluates Partial Embedding Updates with a combination of LoRA, without considering DP-PEFT methods such as DP-Adapter [23, 24, 25, 26], DP-Compacter [27] and DP-BitFit [28] which are often considered in works regarding PEFT and DP-PEFT methods [29, 30, 31].
@@ V-F Large Cohort Noise-Level Simulation
Following [45] and [75], we simulate the noise-level of larger cohort sizes with smaller ones. In practice, differentially private federated learning (DP-FL) is applied to IoT systems with millions of clients [22]. However, it is infeasible to simulate this many clients due to resource constraints. Since a larger cohort size $C$ leads to less noise added for the same privacy guarantee, we simulate realistically large cohort size $C_{\text{large}}$ with smaller cohort size $C_{\text{small}}$. This makes our results more meaningful in terms of deployment of DP-FL in IoT systems. We follow the approach in [45] and [75] for achieving this. The noise $\sigma$ we use for simulation is then computed as $\sigma=\frac{C_{\text{small}}}{C_{\text{large}}}z_{\text{large}}\cdot S$ where $z_{\text{large}}$ is the noise multiplier calculated based on $C_{\text{large}}$.
@@ VI Experiments
We experimentally investigate full fine-tuning, existing parameter-efficient fine-tuning (PEFT) methods and our proposed method DP-DyLoRA under differentially private federated learning (DP-FL) unless otherwise stated. Our experiments cover three different domains, namely, natural language understanding, computer vision and speech in order to ensure that our results are applicable to a wide range of tasks and a variety of scenarios in terms of deployment in IoT systems. We additionally study the impact of data heterogeneity on DP-FL with full fine-tuning to investigate the root causes of the significant performance drop in such learning paradigm.
@@ VI-A Full Fine-tuning
  • When training large models on-device in an IoT system using full fine-tuning under DP-FL, tens of millions of clients may be required for the model to learn effectively.
@@ VI-C Parameter-efficient Fine-tuning
As shown in Figure 6, private models fine-tuned with LoRA significantly outperform those obtained via full fine-tuning on all tasks except for CIFAR-10, where the performance gap between non-private and differentially private training is relatively small. However, on other tasks such as Sentiment140, R8 and Speech Commands, parameter-efficient fine-tuning unlike full fine-tuning allows us to achieve most performance of non-private federated learning with only 1 million clients and a stringent privacy budget of $\epsilon=2$. Moreover, when LoRA is applied, the number of trainable parameters decreases to approximately 1% of that of full fine-tuning. This not only helps reduce computation cost on both the server and client devices but also makes communication between the server and clients much cheaper. Since only the trainable parameters need to be shared, this also serves as an effective solution for potential communication bottleneck when deploying large models in IoT systems.

### v3

@@ Abstract
Federated learning (FL) allows clients to collaboratively train a global model without sharing their local data with a server. However, clients’ contributions to the server can still leak sensitive information. Differential privacy (DP) addresses such leakage by providing formal privacy guarantees, with mechanisms that add randomness to the clients’ contributions. The randomness makes it infeasible to train large transformer-based models, common in modern federated learning systems. In this work, we empirically evaluate the practicality of fine-tuning large scale on-device transformer-based models with differential privacy in a federated learning system. We conduct comprehensive experiments on various system properties for tasks spanning a multitude of domains: speech recognition, computer vision (CV) and natural language understanding (NLU). Our results show that full fine-tuning under differentially private federated learning (DP-FL) generally leads to huge performance degradation which can be alleviated by reducing the dimensionality of contributions through parameter-efficient fine-tuning (PEFT). Our benchmarks of existing DP-PEFT methods show that DP-Low-Rank Adaptation (DP-LoRA) consistently outperforms other methods. An even more promising approach, DyLoRA, which makes the low rank variable, when naively combined with FL would straightforwardly break differential privacy. We therefore propose an adaptation method that can be combined with differential privacy and call it DP-DyLoRA. Finally, we are able to reduce the accuracy degradation and word error rate (WER) increase due to DP to less than 2% and 7% respectively with 1 million clients and a stringent privacy budget of $\bm{\epsilon}=2$.
@@ I Introduction
Today, transformer-based models [1] are becoming increasingly common for a wide range of applications such as natural language understanding (NLU), automatic speech recognition (ASR) and image classification [2, 3]. Compared to models such as recurrent neural networks (RNNs) and convolutional neural networks (CNNs), transformer-based models are known to have several advantages such as being better at handling long-range input dependencies and more efficient for training and inference due to parallel input processing [1]. Pre-training and fine-tuning transformers is the dominant approach for building models with state-of-the-art performance [4, 5]. These models are particularly suitable for deployment on edge devices since they can be pre-trained on massive unlabelled data at the central server without much human effort, and only a small amount of data is required per client when fine-tuning in collaboration with other clients for downstream tasks.
@@
Comprehensive experiments for DP-PEFT methods are missing from the literature. Experiments in existing works [19, 20, 21] fail to address realistic system properties such as a massive number (millions) of clients in a federated learning system [22]. Works such as [20] and [21] show experiments for only a single domain and a single type of DP-PEFT method. Only speech is considered in [19], but it only evaluates Partial Embedding Updates with a combination of LoRA, without considering DP-PEFT methods such as DP-Adapter [23, 24, 25, 26], DP-Compacter [27] and DP-BitFit [28] which are often considered in works regarding PEFT and DP-PEFT methods [29, 30, 31].
@@ V-F Large Cohort Noise-Level Simulation
Following [45] and [75], we simulate the noise-level of larger cohort sizes with smaller ones. In practice, differentially private federated learning (DP-FL) is applied to systems with millions of clients [22]. However, it is infeasible to simulate this many clients due to resource constraints. Since a larger cohort size $C$ leads to less noise added for the same privacy guarantee, we simulate realistically large cohort size $C_{\text{large}}$ with smaller cohort size $C_{\text{small}}$. This makes our results more meaningful in terms of practical deployment of DP-FL. We follow the approach in [45] and [75] for achieving this. The noise $\sigma$ we use for simulation is then computed as $\sigma=\frac{C_{\text{small}}}{C_{\text{large}}}z_{\text{large}}\cdot S$ where $z_{\text{large}}$ is the noise multiplier calculated based on $C_{\text{large}}$.
@@ VI Experiments
We experimentally investigate full fine-tuning, existing parameter-efficient fine-tuning (PEFT) methods and our proposed method DP-DyLoRA under differentially private federated learning (DP-FL) unless otherwise stated. Our experiments cover three different domains, namely, natural language understanding, computer vision and speech in order to ensure that our results are applicable to a wide range of tasks and a variety of scenarios. We additionally study the impact of data heterogeneity on DP-FL with full fine-tuning to investigate the root causes of the significant performance drop in such learning paradigm.
@@ VI-A Full Fine-tuning
  • When training large models on-device using full fine-tuning under DP-FL, tens of millions of clients may be required for the model to learn effectively.
@@ VI-C Parameter-efficient Fine-tuning
As shown in Figure 6, private models fine-tuned with LoRA significantly outperform those obtained via full fine-tuning on all tasks except for CIFAR-10, where the performance gap between non-private and differentially private training is relatively small. However, on other tasks such as Sentiment140, R8 and Speech Commands, parameter-efficient fine-tuning unlike full fine-tuning allows us to achieve most performance of non-private federated learning with only 1 million clients and a stringent privacy budget of $\epsilon=2$. Moreover, when LoRA is applied, the number of trainable parameters decreases to approximately 1% of that of full fine-tuning. This not only helps reduce computation cost on both the server and client devices but also makes communication between the server and clients much cheaper. Since only the trainable parameters need to be shared, this also serves as an effective solution for potential communication bottleneck when deploying large models on edge devices.